# Logarithmic critical slowing down in complex systems: from statics to dynamics

### v1

@@ I Introduction
holding at the dynamical $A_{2}$ singularity of the MCT for undercooled viscous liquids. Changing the values of the external parameters a dynamic arrest line can be drawn in the phase diagram consisting of $A_{2}$ points. In given systems the exponent parameter $\lambda$ tends to $1$ along the dynamic arrest line, approaching a $A_{3}$ point. In that limit the exponent $a$ tends to zero and logarithmic corrections become relevant to the relaxation.
Hereafter, we present a general method for the quantitative computation of the coefficient of the logarithmic decay of the density-density correlation functions in viscous liquids. Göetze and Sjögren Gotze and Sjogren (1989) predicted a $1/\ln t^{2}$ decay exactly at the $A_{3}$ singularity and a behavior of the kind $-(\ln t)^{\gamma}$ as $a\sim 0$, $\lambda\sim 1$. Their exemplifying case is the $F_{13}$ mode-coupling schematic theory. However, in this case, the $A_{3}$ singularity cannot be directly accessed in experiments or in numerical simulations because it occurs in the region of the phase diagram pertaining to the glassy phase, beyond the dynamic arrest line where TTI breaks down and the MCT does not hold anymore. Therefore, in the liquid phase the presence of the $A_{3}$ singularity is only felt in weakly logarithmic corrections to the power-law $\beta$ decay in regions of the space parameters (temperature, packing fraction, …) close enough to it. Other systems displaying this kind of singularity include disordered spin-glass models Crisanti et al. (1993a); Crisanti and Leuzzi (2006, 2015); Franz et al. (2013), liquids in porous media both in the MCT Krakoviack (2005, 2007) and in the HNC approximations Madden and Glandt (1988); Given and Stell (1992) and liquid models with pinned particles Cammarota and Biroli (2012). The $-\ln t$ behavior of the correlation function appears to be the correct fitting law for about a decade or two in most of the known experiments and numerical simulations of repulsive colloids Sjogren (1991); Zaccarelli et al. (2002); Sciortino et al. (2003a, b); Götze (2009). Also cases where the $A_{3}$ singularity is directly accessible from the liquid phase are devised in MCT, for instance in the $F_{12}$ schematic model. The quantitative estimation of the parameters of the logarithmic behavior can, so far, be performed exactly only in MCT schematic theories. Moreover, as, e.g., in the case of the $F_{13}$ model, the $A_{3}$ singularities lie in the region where one of the fundamental assumptions on which MCT is built, TTI, does not hold.
@@ II Outline of the Results
$Q_{ab}={1\over N}\sum_{i=1}^{N}\overline{\langle s_{i}^{a}s_{i}^{b}\rangle}\ .$ (2)
For the case $n\to 0$ the angle brackets are thermal averages and the overline is the average over the quenched disorder. The case $n\to 1$ applies to problems where for each disorder realization there are many metastable excited states, whose number grows exponentially with the size of the system. In this latter case, then, the angle brackets represent thermal averages inside a metastable state and the overline represents averages over different metastable states and over the quenched disorder.
@@
and $\delta\rho_{a}(\bm{k})$ is the fluctuation of $\rho_{a}(\bm{k})$ with respect to its average $\langle\rho_{a}(\bm{k})\rangle$. We note that choice of $\bm{k}$ in Eq. (3) is free arbitrary and one could consider, instead, the mean-square-displacement Parisi et al. (2020). We refer the reader to section II.B of Rizzo (2016) for a thorough discussion on the choice of the order parameter.
@@
where the $J^{\prime}s$ are quenched random interactions and the $s_{i}$ can be either Ising spin or satisfy a spherical constraint. In the spherical $3$-spin case in presence of a magnetic field there is a tricritical point with $\lambda=1$ in the temperature-magnetic field plane where a line of discontinuous transitions meets a line of continuous transitions Crisanti and Sommers (1992); Crisanti et al. (1993a). Another example is that of a mixed $2+3$ model that corresponds to the so-called schematic $F_{12}$ model in the context of MCT. In the phase diagram, e.g., in the plane of the magnitudes of the $2$-spin and the $3$-spin interaction there is a $\lambda=1$ critical point. Upon increasing the relative magnitude of the $3$-spin interaction, a line of continuous transitions meets a line of discontinuous transitions Franz et al. (2013); Caltagirone et al. (2013). Random pinning of a spherical $p$-spin glass model, i.e. freezing a fraction $c$ of the spins Cammarota and Biroli (2012), is also relevant: in the temperature-concentration plane there is a line of discontinuous transitions that, upon increasing the concentration, ends in a point characterized by $\lambda=1$. Finally we mention the Potts Spin-Glass with Hamiltonian:
@@ III.1 Theories with $n=0$
$=$ $\left\{\begin{array}[]{c c}(\bar{\theta}_{1}-\bar{\theta}_{2})\theta_{2}&,\quad t_{1}>t_{2}\\ &\\ (\bar{\theta}_{1}-\bar{\theta}_{2})\theta_{1}&,\quad t_{2}>t_{1}.\end{array}\right.$

### v3

@@ I Introduction
holding at the dynamical $A_{2}$ singularity of the MCT for undercooled viscous liquids, $\Gamma(x)$ being the gamma function. Changing the values of the external parameters a dynamic arrest line can be drawn in the phase diagram consisting of $A_{2}$ points. In given systems the exponent parameter $\lambda$ tends to $1$ along the dynamic arrest line, approaching a $A_{3}$ point. In that limit the exponent $a$ tends to zero and logarithmic corrections become relevant to the relaxation.
Hereafter, we present a general method for the quantitative computation of the coefficient of the logarithmic decay of the density-density correlation functions in viscous liquids. Göetze and Sjögren Gotze and Sjogren (1989) predicted a $1/\ln t^{2}$ decay exactly at the $A_{3}$ singularity and a behavior of the kind $-(\ln t)^{\gamma}$ as $a\sim 0$, $\lambda\sim 1$. Their exemplifying case is the $F_{13}$ mode-coupling schematic theory Götze (2009). However, in this case, the $A_{3}$ singularity cannot be directly accessed in experiments or in numerical simulations because it occurs in the region of the phase diagram pertaining to the glassy phase, beyond the dynamic arrest line where TTI breaks down and the MCT does not hold anymore. Therefore, in the liquid phase the presence of the $A_{3}$ singularity is only felt in weakly logarithmic corrections to the power-law $\beta$ decay in regions of the space parameters (temperature, packing fraction, …) close enough to it. Other systems displaying this kind of singularity include disordered spin-glass models Crisanti et al. (1993a); Crisanti and Leuzzi (2006, 2015); Franz et al. (2013); Cammarota and Biroli (2012a), liquids in porous media both in the MCT Krakoviack (2005, 2007) and in the hypernetted-chain approximations Madden and Glandt (1988); Given and Stell (1992) and liquid models with pinned particles Cammarota and Biroli (2012b). The $-\ln t$ behavior of the correlation function appears to be the correct fitting law for about a decade or two in most of the known experiments and numerical simulations of repulsive colloids Sjogren (1991); Zaccarelli et al. (2002); Sciortino et al. (2003a, b); Götze (2009). Also cases where the $A_{3}$ singularity is directly accessible from the liquid phase are devised in MCT, for instance in the $F_{12}$ schematic model Götze (2009). The quantitative estimation of the parameters of the logarithmic behavior can, so far, be performed exactly only in MCT schematic theories. Moreover, as, e.g., in the case of the $F_{13}$ model, the $A_{3}$ singularities lie in the region where one of the fundamental assumptions on which MCT is built, TTI, does not hold.
@@ II Outline of the Results
$Q_{ab}={1\over N}\sum_{i=1}^{N}\overline{\langle s_{i}^{a}s_{i}^{b}\rangle}$ (2)
where the $s_{i}$, $i=1,\dots,N$, are spins. For the case $n\to 0$ the angle brackets are thermal averages and the overline is the average over the quenched disorder. The case $n\to 1$ applies to problems where for each disorder realization there are many metastable excited states, whose number grows exponentially with the size of the system. In this latter case, then, the angle brackets represent thermal averages inside a metastable state and the overline represents averages over different metastable states and over the quenched disorder.
@@
and $\delta\rho_{a}(\bm{k})$ is the fluctuation of $\rho_{a}(\bm{k})$ with respect to its average $\langle\rho_{a}(\bm{k})\rangle$. We note that choice of $\bm{k}$ in Eq. (3) is arbitrary and one could consider, instead, the mean-square-displacement Parisi et al. (2020). We refer the reader to section II.B of Rizzo (2016) for a thorough discussion on the choice of the order parameter.
@@
where the $J^{\prime}s$ are quenched random interactions and the $s_{i}$ can be either Ising spin or satisfy a spherical constraint. In the spherical $3$-spin case in presence of a magnetic field there is a tricritical point with $\lambda=1$ in the temperature-magnetic field plane where a line of discontinuous transitions meets a line of continuous transitions Crisanti and Sommers (1992); Crisanti et al. (1993a). Another example is that of a mixed $2+3$ model that corresponds to the so-called schematic $F_{12}$ model in the context of MCT. In the phase diagram, e.g., in the plane of the magnitudes of the $2$-spin and the $3$-spin interaction there is a $\lambda=1$ critical point. Upon increasing the relative magnitude of the $3$-spin interaction, a line of continuous transitions meets a line of discontinuous transitions Franz et al. (2013); Caltagirone et al. (2013). Random pinning of a spherical $p$-spin glass model, i.e. freezing a fraction $c$ of the spins Cammarota and Biroli (2012b, a), is also relevant: in the temperature-concentration plane there is a line of discontinuous transitions that, upon increasing the concentration, ends in a point characterized by $\lambda=1$. Finally we mention the Potts Spin-Glass with Hamiltonian:
@@ III.1 Theories with $n=0$
$=$ $\left\{\begin{array}[]{c c}(\bar{\theta}_{1}-\bar{\theta}_{2})\theta_{2}&,\quad t_{1}>t_{2}\\ &\vspace*{-.2cm}\hfil\\ (\bar{\theta}_{1}-\bar{\theta}_{2})\theta_{1}&,\quad t_{2}>t_{1}.\end{array}\right.$
@@ VI Conclusions
It is also interesting to mention two instances in which there is instead no connection between logarithmic slowing down and connected correlation functions of the overlap fluctuations. This is provided by Fredrikson-Andersen Kinetically constrained model on the Bethe lattice with either random pinning Ikeda et al. (2017) or with non-homogeneous facilitation Sellitto et al. (2010); Arenzon and Sellitto (2012): the analytical solution of these models Perrupato and Rizzo (2022) has indeed allowed to demonstrate and quantify the logarithmic slowing down as given by eq. (16) but a thermodynamic analysis of the model has revealed that there is no connection between the observed MCT-like dynamics and connected correlation functions of the overlap that are not divergent at the critical point Perrupato and Rizzo (2023), i.e. eq. (18) does not hold.
The results presented here have a mean-field nature and their relevance for realistic models in finite dimension, say two and three, in principle is not granted. This issue notwithstanding, analytical predictions like those derived here and more generally those of Mode-Coupling-Theory are typically used to describe successfully numerical and experimental data in the context of supercooled liquids and colloids Götze (2009); Sjogren (1991); Zaccarelli et al. (2002); Sciortino et al. (2003a, b). A relevant exception is the discontinuous dynamical arrest transition that is known to become a cross-over in finite dimension: it has been argued that this phenomenon is precisely due to long-wavelength fluctuations that destroy the mean-field theory below the upper critical dimension eight Franz et al. (2011); Rizzo (2014, 2016); Rizzo and Voigtmann (2015, 2020); Elizondo-Aguilera et al. (2022). In the Spin-Glass literature the relevance of mean-field theory for realistic systems is an essential question Bray and Roberts (1980); Lubensky et al. (2023); Angelini et al. (2022) but it is typically discussed in a purely static context and there is at present no understanding of the fate of the dynamical power-law (and logarithmic) decays found in mean-field theory.
It should be also noted that we have been considering equilibrium dynamics, therefore the results are limited to the temperature regime in which the system can still be thermalized and there is no aging. Nonetheless it is not unlikely that, with more effort, the analysis can be extended to the aging regime, as it is known that static replicated theories can be connected to off-equilibrium dynamics as well Cugliandolo and Kurchan (1993); Caltagirone et al. (2012d).